10.6 A 1.00 m^3 element of soil is located below the groundwater table. When a new...

10.6

A 1.00 m^3 element of soil is located below the groundwater table. When a new Compressive load was applied, this element consolidated, producing a vertical strain, ℇz, of 8.5%. The horizontal strain was zero. Compute the volume of water squeezed out of this soil during consolidation and express your answer in liters.
Answer #1

Assume the size of soil 1m*1m*1m

Initial volume = 1m3

Vertical strain = 8.5%

Change in depth = (8.5/100)*1

= 0.085m

Change in volume = 1*1*0.085

= 0.085m3​​​​​​

This change is due to consolidation(expulsion of water from pores of soil)

Amount of water expelled = 0.085*1000

= 85 litres